OpenStudy (anonymous):

The area of a trapezoid is 144 ft2. If the height is 8 feet and one of the bases is 12 feet, what is the length of the other base?

Directrix (directrix):

Other base is 24. Work to follow.

Directrix (directrix):

A = (1/2) h (b1 + b2) where b-one and b-two are the bases. 144 = (1/2) (8) (12 + b) 144 = 4 ( 12 + b) 144 = 48 + 4b 96 = 4b 24 = b

Directrix (directrix):

Check my work, please. :)

OpenStudy (anonymous):

Your correct. Ty :D